OpenStudy (mitu12):

a student got 44 questions correct on his test with 55 questions. what percent was NOT correct

OpenStudy (tester97):

\[55-44=11\] \[11~not~correct\] Make that into a fraction \[\frac{ 11 }{ 55 }\] simplify. \[\frac{ 1 }{ 5 }\] convert to percentage 20%

OpenStudy (mitu12):

thank you so much

OpenStudy (tester97):

welcome :)